### v2

# ERROR ESTIMATION FOR FINITE ELEMENT SOLUTIONS ON MESHES THAT CONTAIN THIN ELEMENTS

Abstract.
In an error estimation of finite element solutions to the Poisson equation, we usually impose the shape regularity assumption on the meshes to be used. In this paper, we show that even if the shape regularity condition is violated, the standard error estimation can be obtained if “bad” elements elements that violate the shape regularity or maximum angle condition are covered virtually by simplices that satisfy the minimum angle condition. A numerical experiment illustrates the theoretical result.
Keywords: finite element method, triangulation, minimum and maximum angle condition, shape regularity condition, bad triangles
MSC 2020: 65D05, 65N30
00footnotetext: The authors were supported by JSPS KAKENHI Grant Numbers 24K00538 and 21K03372.
1. Introduction
The mathematical theory of finite element methods has been well established, and many well written textbooks on the subject are available [5, 7, 10]. In these textbooks, the central assumption of the theory is the shape regularity condition on the meshes that are used to define finite element solutions. The shape regularity condition requires that each element is not very degenerate. For example, if a mesh consists of triangles, each triangle should not be very “flat”. It is known that the shape regularity condition is equivalent to the minimum angle condition.
However, suppose the problem domain has a highly complex shape, and/or the problem has an anisotropic nature, and we introduce an adaptive mesh refinement procedure. In that case, the meshes might contain some very thin elements and the error estimation of the finite element method may thus lose its theoretical validity.
There are two ways to overcome this issue. One way is to introduce a new geometric condition to meshes, where the condition ensures a standard convergence rate of finite element solutions. It is well known that if the maximum angles of triangles in a mesh are bounded by a fixed constant $\theta<\pi$, then the standard error estimation of finite element solutions holds [2]. This condition is known as the maximum angle condition. The approach has been extended in many ways. See, for example, [1, 12, 13, 15].
The other way was initiated by Kučera [17] and Duprez–Lleras–Lozinski [8]. Suppose that a mesh contains some “bad” elements. In [17, 8], if bad elements satisfy some conditions, the usual Lagrange interpolation is modified in certain ways on bad elements so that the standard error estimation holds.
This paper extends their approach. We show that if all bad elements are covered virtually by simplices that satisfy the minimum angle condition, the standard error estimation holds. Note that the simplices are virtual and do not need to be elements of the mesh considered.
In Section 2, we present the model problem, the precise definitions of conformal meshes, and the above-mentioned conditions. In addition, we give an example of bad meshes for which the standard error estimation holds. Section 3 presents detailed assumptions of meshes and the main theorem. Section 4 presents a simpler version of the assumption and the main theorem for the case $n=2$. In Section 5, we compare our approach with approaches adopted in prior works [17, 8]. In Section 6, we present a numerical experiment that confirms the theoretical result.
2. Preliminary
2.1. Model problem
Let $\Omega\subset\mathbb{R}^{n}$ $(n=2,3)$ be a bounded polygonal or polyhedral domain. For simplicity, we consider the Poisson equation
$-\Delta u=f\;\text{ in }\;\Omega,\qquad u=0\;\text{ on }\;\partial\Omega$
and its weak form
(2.1) $\int_{\Omega}\nabla u\cdot\nabla v\,\mathrm{d}\mathbf{x}=\int_{\Omega}fv\,\mathrm{d}\mathbf{x}\quad\forall v\in H_{0}^{1}(\Omega)$
as the model problem, where $f\in L^{2}(\Omega)$ is a given function. Here, $L^{p}(\Omega)$, $H^{m}(\Omega)$, $H_{0}^{1}(\Omega)$, and $W^{m,p}(\Omega)$ ($m=1,2$, $1\leq p\leq\infty$) are the standard Lebesgue and Sobolev spaces with norms and seminorms $\|\cdot\|_{L^{p}(\Omega)}$, $\|\cdot\|_{H^{m}(\Omega)}$, $|\cdot|_{H^{m}(\Omega)}$, $\|\cdot\|_{W^{m,p}(\Omega)}$, and $|\cdot|_{W^{m,p}(\Omega)}$. In this paper, we assume that the exact solution $u$ of (2.1) belongs to $H^{2}(\Omega)$.
2.2. Finite element method
We approximate the exact solution of (2.1) using the standard piecewise linear finite element method. To this end, we define the piecewise linear finite element spaces by
$S_{h}:=\left\{v_{h}\in C^{0}(\overline{\Omega})\bigm{|}v_{h}|_{\tau}\in\mathcal{P}_{1},\forall\tau\in\mathcal{T}_{h}\right\},\quad S_{h0}:=S_{h}\cap H_{0}^{1}(\Omega),$
where $\mathcal{P}_{1}$ is the set of polynomials with degree at most $1$. Here, $\mathcal{T}_{h}$ is a conformal simplicial mesh of $\Omega$. That is, $\mathcal{T}_{h}$ is a finite set of $n$-simplices that satisfies the conditions
  • $\bigcup_{\tau\in\mathcal{T}_{h}}\tau=\overline{\Omega}$.
  • If $\tau_{1}\cap\tau_{2}\neq\emptyset$ for $\tau_{1}$, $\tau_{2}\in\mathcal{T}_{h}$ with $\tau_{1}\neq\tau_{2}$, then $\tau_{1}\cap\tau_{2}$ is their common $r$-face ($r<n$) (that is, there are no hanging nodes in $\mathcal{T}_{h}$).
As usual, we define $h_{\tau}:=\mathrm{diam}\,\tau$ and $h:=\max_{\tau\in\mathcal{T}_{h}}h_{\tau}$. The finite element solution $u_{h}\in S_{h0}$ is then defined as the unique solution of the equation
(2.2) $\int_{\Omega}\nabla u_{h}\cdot\nabla v_{h}\,\mathrm{d}\mathbf{x}=\int_{\Omega}fv_{h}\,\mathrm{d}\mathbf{x},\quad\forall v_{h}\in S_{h0}.$
2.3. Céa’s lemma and interpolations
For the error of $u_{h}$, Céa’s lemma claims that there exists a constant $C>0$ such that
$|u-u_{h}|_{H^{1}(\Omega)}\leq{C}\inf_{w_{h}\in S_{h0}}|u-w_{h}|_{H^{1}(\Omega)}\leq{C}\,|u-\Pi^{*}u|_{H^{1}(\Omega)},$
where $\Pi^{*}u\in S_{h}$ is any interpolation of the exact solution $u$. Therefore, if we have
(2.3) $|u-\Pi^{*}u|_{H^{1}(\Omega)}\leq Ch|u|_{H^{2}(\Omega)},$
we obtain the standard convergence rate
$|u-u_{h}|_{H^{1}(\Omega)}\leq Ch|u|_{H^{2}(\Omega)}.$
Usually, the Lagrange interpolation operator is used as $\Pi^{*}$. The piecewise linear Lagrange interpolation $\Pi_{\tau}^{1}u\in\mathcal{P}_{1}$ is defined by
$(\Pi_{\tau}^{1}u)(\mathbf{x}_{i})$ $=u(\mathbf{x}_{i})\text{ on $n$-simplex }\tau\text{ with vertices}\mathbf{x}_{i},i=1,\dots,n+1,$
$\Pi_{h}^{1}u\bigm{|}_{\tau}$ $=\Pi_{\tau}^{1}u,\qquad\forall\tau\in\mathcal{T}_{h}.$
In the next subsection, we mention conditions for which the Lagrange interpolation satisfies the standard error estimation
(2.4) $|u-\Pi_{\tau}^{1}u|_{W^{1,p}(\tau)}\leq C_{p}h_{\tau}|u|_{W^{2,p}(\tau)},\quad\forall u\in W^{2,p}(\tau).$
2.4. Conditions and constants related to meshes
Definition 1.
For an $n$-simplex $\tau$, $\theta_{\tau}^{min}$ and $\theta_{\tau}^{max}$ denote the minimum and maximum inner angles of $\tau$, respectively. Here, the inner angles of a tetrahedron mean, for $\theta_{\tau}^{min}$, the inner angles of each face and the solid angles at vertices, and for $\theta_{\tau}^{max}$, the inner angles of each face and the dihedral angles [3, 16].
The following conditions are well known for geometric conditions of simplices.
Definition 2.
$(1)$ If $\theta_{\tau}^{min}\geq\alpha>0$, $\tau$ is said to satisfy the minimum angle condition with respect to $\alpha$.
$(2)$ If $\theta_{\tau}^{max}\leq\beta<\pi$, $\tau$ is said to satisfy the maximum angle condition with respect to $\beta$.
If $\tau$ satisfies the minimum angle condition with respect to $\alpha$, the error estimation (2.4) holds for any $p$, $1\leq p\leq\infty$ with $C_{p}=C_{\alpha,p}^{min}$. Moreover, there exists a constant $D_{\alpha}$ such that
(2.5) $\frac{h_{\tau}}{\rho_{\tau}}\leq D_{\alpha},$
where $\rho_{\tau}$ is the diameter of the inscribed ball in $\tau$. This condition is known as the shape regularity condition with respect to $D_{\alpha}$. It is well known that the shape regularity condition is equivalent to the minimum angle condition.
Suppose that $\tau$ satisfies the maximum angle condition with respect to $\beta$. If $n=2$, (2.4) holds for any $p$, $1\leq p\leq\infty$ with $C_{p}=C_{\beta,p}^{max}$. If $n=3$, (2.4) holds for $p$, $2<p\leq\infty$ with $C_{p}=C_{\beta,p}^{max}$; however, $C_{\beta,p}^{max}\to\infty$ as $p\searrow 2$, see [9, 14, 16, 18].
To specify the characteristics of meshes, we introduce constants $\theta$, $\psi$, where $30<\theta\leq\pi/$ 3, $30<\psi\leq\pi/$ 3, and fix them throughout this paper. We define
$K_{\theta}$ :=$:=\begin{cases}\;\{\tau\subset\mathbb{R}^{2}\mid\tau\text{ is a triangle and }\theta_{\tau}^{max}\leq\pi-{2}\theta\},&\text{for $n=2$},\\ \;\{\tau\subset\mathbb{R}^{3}\mid\tau\text{ is a tetrahedron and }\theta_{\tau}^{min}\geq\theta\},&\text{for $n=3$}.\end{cases}$
A simplex $\tau$ that does not belong to $K_{\theta}$ is regarded as “bad”, a simplex $\tau$ that satisfies the minimum angle condition $\theta_{\tau}^{min}\geq\theta$ is regarded as “good”, and other types of simplex are regarded as “ordinary”. Note that for $n=2$, $K_{\theta}$ consists of “good” and “ordinary” triangles, and for $n=3$, all tetrahedrons in $K_{\theta}$ are “good”.
For a simplex $\tau$ that satisfies $\tau\in K_{\theta}$ or $\theta_{\tau}^{min}\geq\psi$, we rewrite (2.4) as
(2.6) $|u-\Pi_{\tau}^{1}u|_{H^{1}(\tau)}\leq A_{\theta,\psi}h_{\tau}|u|_{H^{2}(\tau)},\qquad\forall u\in H^{2}(\tau),$
where $A_{\theta,\psi}$ is a constant that depends only on $\theta$ and $\psi$. The existence of $A_{\theta,\psi}$ is assured by the minimum and maximum angle conditions. (The parameter $\theta$ is used to categorize elements in mesh whereas the parameter $\psi$ relates to virtual simplices that appear later.)
2.5. Sobolev embedding theorem
Let $\hat{\tau}$ be a reference simplex with $h_{\hat{\tau}}=1$. Then, by the Sobolev embedding theorem [6], we have
$H^{2}(\hat{\tau})\subset W^{1,6}(\hat{\tau})\subset L^{\infty}(\hat{\tau})$
and
$\|\hat{u}-\Pi_{\hat{\tau}}^{1}\hat{u}\|_{L^{\infty}(\hat{\tau})}\leq\widehat{B}|\hat{u}-\Pi_{\hat{\tau}}^{1}\hat{u}|_{H^{2}(\hat{\tau})}=\widehat{B}|\hat{u}|_{H^{2}(\hat{\tau})},\quad\hat{u}\in H^{2}(\hat{\tau}),$
where $\widehat{B}$ is a constant that depends only on $\hat{\tau}$. Here, we used the fact that $(u-\Pi_{\hat{\tau}}u)(\mathbf{x}_{i})=0$, $i=1,\dots,n+1$, and $\|\hat{u}-\Pi_{\hat{\tau}}^{1}\hat{u}\|_{H^{2}(\hat{\tau})}$ may be bounded by $|\hat{u}-\Pi_{\hat{\tau}}^{1}\hat{u}|_{H^{2}(\hat{\tau})}$ up to a constant. For a general simplex $\tau$ with $\theta_{\tau}^{min}\geq\psi$, we consider an affine map $\eta:\hat{\tau}\to\tau$, $\eta(\mathbf{x})=A\mathbf{x}+\mathbf{b}$, where $A$ is an $n\times n$ nonsingular matrix and $\mathbf{b}\in\mathbb{R}^{n}$. Each $u\in H^{2}(\tau)$ is pulled back to $\hat{u}$ as $\hat{u}(\mathbf{x}):=u(\eta(\mathbf{x}))$. We then have
$\|u-\Pi_{\tau}^{1}u\|_{L^{\infty}(\tau)}^{2}$ $=\|\hat{u}-\Pi_{\hat{\tau}}^{1}\hat{u}\|_{L^{\infty}(\hat{\tau})}^{2}\leq\widehat{B}^{2}|\hat{u}|_{H^{2}(\hat{u})}^{2}$
$\leq\widehat{B}^{2}|\det A|^{-1}\|A\|^{4}|u|_{H^{2}(\tau)}^{2},$
where $\|A\|$ is the matrix norm of $A$ with respect to the Euclidean norm. We denote the Lebesgue measure of $\tau$ by $|\tau|$. By [7], the coefficient of the last term is estimated as
$\det A=\frac{|\tau|}{|\hat{\tau}|}\geq Ch^{n},\qquad\|A\|\leq\frac{h_{\tau}}{\rho_{\hat{\tau}}},$
and we finally obtain the embedding inequality
(2.7) $\|u-\Pi_{\tau}^{1}u\|_{L^{\infty}(\tau)}\leq B_{\psi}h_{\tau}^{2-n/2}|u|_{H^{2}(\tau)},\quad\forall u\in H^{2}(\tau),$
where $B_{\psi}$ is a constant that depends on $\psi$.
2.6. Example of error estimation on a bad mesh.
The minimum or maximum angle conditions are sufficient but not necessary for the standard error estimation (2.4) or (2.6). Suppose that we have the two meshes $\mathcal{T}_{h}^{1}$ and $\mathcal{T}_{h}^{2}$ depicted in Figure 1 (where $\mathcal{T}_{h}^{2}$ has already been mentioned in [11]). The mesh $\mathcal{T}_{h}^{1}$ consists of congruent isosceles right triangles and satisfies the minimum angle (shape regularity) condition. In contrast, $\mathcal{T}_{h}^{2}$ is a “refinement” of $\mathcal{T}_{h}^{1}$ and contains very flat triangles. Thus, $\mathcal{T}_{h}^{2}$ does not satisfy either the minimum or maximum angle condition when the flat triangles degenerate.
Figure  1. Mesh $\mathcal{T}_{h}^{1}$ (left) and its refinement $\mathcal{T}_{h}^{2}$ (right).
As each $\tau\in\mathcal{T}_{h}^{2}$ is a subset of some $\tau^{\prime}\in\mathcal{T}_{h}^{1}$, we define the interpolation $\Pi^{*}$ on $\mathcal{T}_{h}^{2}$ as follows. Let $\mathcal{N}_{h}^{1}$ and $\mathcal{N}_{h}^{2}$ be the sets of vertices of $\mathcal{T}_{h}^{1}$ and $\mathcal{T}_{h}^{2}$, respectively. Let $\{\varphi_{i}\}\subset S_{h}$ be the basis of $S_{h}$ on $\mathcal{T}_{h}^{2}$ defined by
$\varphi_{i}(\mathbf{x}_{i})=1,\qquad\varphi_{i}(\mathbf{x}_{j})=0,\quad\mathbf{x}_{i},\mathbf{x}_{j}\in\mathcal{N}_{h}^{2},\quad i\neq j.$
We define
$w_{i}:=\begin{cases}(\Pi_{\mathcal{T}_{h}^{1}}^{1}u)(\mathbf{x}_{i}),&\text{for }\;\mathbf{x}_{i}\in\mathcal{N}_{h}^{2}\backslash\mathcal{N}_{h}^{1},\\ u(\mathbf{x}_{i}),&\mathbf{x}_{i}\in\mathcal{N}_{h}^{1},\end{cases}\qquad\Pi^{*}u:=\sum_{i}w_{i}\varphi_{i},$
where $\Pi_{\mathcal{T}_{h}^{1}}^{1}$ is the Lagrange interpolation on $\mathcal{T}_{h}^{1}$. As $\mathcal{T}_{h}^{1}$ is a good mesh, the standard error estimation (2.6) holds, and the finite element solutions converge to $u$ with $\mathcal{O}(h)$. Moreover, $\Pi^{*}$ inherits these properties because $\Pi_{\mathcal{T}_{h}^{1}}^{1}u=\Pi^{*}u$ on $\Omega$.
This example reminds us that even if a mesh contains bad elements, we might be able to define an interpolation $\Pi^{*}$ that satisfies (2.3) if bad elements are covered by fine simplices under some additional conditions.
3. Assumptions and the main theorem
In this section, we introduce assumptions that ensure the standard error estimation (2.3) on meshes that contain bad elements and explain the main result in detail. Recall that $30<\theta,\psi\leq\pi/$ 3 are given fixed constants. Let $M$, $N$ be positive integers and $C\geq 1$. These constants are also fixed throughout the paper.
For a simplex $\tau$, let $v(\tau)$ be the set of vertices of $\tau$. For $n$-simplices ${T_{1}}$ and ${T_{2}}$, we define
$V({T_{1}},{T_{2}}):=\mathrm{Conv}(v({T_{1}})\cap v({T_{2}})),\qquad V_{\partial\Omega}({T_{1}}):=\mathrm{Conv}(v({T_{1}})\cap\partial\Omega),$
where Conv$(S)$ denotes the convex hull of the point set $S$. We present several examples in Figures 2 and 3. For triangles ${T_{1}}$ and ${T_{2}}$, $V({T_{1}},{T_{2}})$ and $V_{\partial\Omega}({T_{1}})$ are drawn with thick lines and dots.
[…]
Figure  2. Examples of $V({T_{1}},{T_{2}})$.
${T_{1}}$$\partial\Omega$
${T_{1}}$$\partial\Omega$
$\partial\Omega$${T_{1}}$
Figure  3. Examples of $V_{\partial\Omega}({T_{1}})$.
We impose the following conditions for the mesh $\mathcal{T}_{h}$.
Assumption 3.
$(1)$ There exist $n$-simplices $T_{1},\dots,T_{m}$ such that $\theta_{T_{k}}^{min}\geq\psi$ and $h_{T_{k}}\leq Ch$ for $k=1,\dots,m$. We assume that $\ring{T}_{k}\subset\Omega$, where $\ring{T}_{k}$ is the interior set of $T_{k}$. The interior sets $\{\ring{T}_{k}\}$ may overlap, and the multiplicity of the overlap of $\{\ring{T}_{k}\}_{k=1}^{m}$ is at most $M$. (Note that we do NOT assume $T_{k}\in\mathcal{T}_{h}$ ($k=1,\dots,m)$.)
$(2)$ The mesh $\mathcal{T}_{h}$ is expressed as a disjoint union $\mathcal{T}_{h}=\bigcup_{k=0}^{m}Q_{k}$ such that $Q_{0}\subset\mathcal{T}_{h}\cap K_{\theta}$ and each “bad” element $\tau\in\mathcal{T}_{h}\backslash K_{\theta}$ belongs to one of $\{Q_{k}\}_{k=1}^{m}$. Furthermore, we assume that
$\bigcup_{\tau\in Q_{k}}\tau\subset\overline{T}_{k},\quad\text{ and }\quad\#Q_{k}\leq N\quad\text{ for }k=1,\dots,m{,}$
where #$\#Q_{k}$ is the number of elements in $Q_{k}$. Note that $Q_{0}$ consists of “good” and “ordinary” elements, and $Q_{k}$ $(k=1,\dots,m)$ may contain “good” or “ordinary” elements as well as “bad” elements.
$(3)$ If $(v(\tau_{0})\cap v(\tau_{1}))\backslash v(T_{k})\neq\emptyset$ for $\tau_{0}\in Q_{0}$, $\tau_{1}\in Q_{k}$, $1\leq k\leq m$, then we assume that $\theta_{\tau_{0}}^{min}\geq{\theta}$. (Note that if $n=3$, this condition is automatically satisfied because all elements belong to $Q_{0}$ are “good”.)
$(4)$ If $v(\tau_{1})\cap v(\tau_{2})\neq\emptyset$ for $\tau_{1}\in Q_{k}$, $\tau_{2}\in Q_{\ell}$, $1\leq k<\ell\leq m$, then we assume $v(\tau_{1})\cap v(\tau_{2})\subset V(T_{k},T_{\ell})$.
$(5)$ If $v(\tau_{1})\cap\partial\Omega\neq\emptyset$ for $\tau_{1}\in Q_{k}$, $1\leq k\leq m$, then we assume $v(\tau_{1})\cap\partial\Omega\subset V_{\partial\Omega}(T_{k})$.
Let $\{\varphi_{i}\}\subset S_{h}$ be the basis of $S_{h}$ on the mesh $\mathcal{T}_{h}$.
Definition 4.
For $u\in H^{2}(\Omega)$, we define the interpolation $\Pi^{*}u$ on $\mathcal{T}_{h}$ by
$w_{i}$ :=$:=\begin{cases}(\Pi_{T_{k}}^{1}u)(\mathbf{x}_{i}),&\text{if $\mathbf{x}_{i}\in v(\tau)$ for some $\tau\in Q_{k}$}\\ u(\mathbf{x}_{i}),&\text{otherwise}\end{cases}$
$\Pi^{*}u$ $:=\sum_{i}w_{i}\varphi_{i},$
where $\mathbf{x}_{i}$ are the vertices of the mesh $\mathcal{T}_{h}$ and $\Pi_{T_{k}}^{1}$ is the Lagrange interpolation on $T_{k}$.
Remark 5. Assumption 3 (1), (2) requires that each bad element in $\mathcal{T}_{h}$ belongs to one of the disjoint element clusters $Q_{k}$ ($k=1,\dots,m$) and each $Q_{k}$ is virtually covered by a simplex $T_{k}$ ($k=1,\dots,m$). The interpolation $\Pi^{*}u$ is defined by $\Pi_{T_{k}}^{1}u$ on each $T_{k}$.
To explain the meaning of Assumption 3 (3), we consider the situation that $\tau_{0}\in Q_{0}$ and $\tau_{1}\in Q_{k}$ as is depicted in Figure 4. Then, $\tau_{1}$ is virtually covered by $T_{k}$ (with other elements in $Q_{k}$, if $\{\tau_{1}\}\subsetneqq Q_{k}$). Let $\mathbf{x}\in v(\tau_{0})\cap v(\tau_{1})\backslash v(T_{k})$. Figure 4 (left) depicts such a situation. As the interpolation $\Pi^{*}u$ is defined by $\Pi_{T_{k}}^{1}u$, we have $(\Pi^{*}u)(\mathbf{x})=(\Pi_{T_{k}}^{1}u)(\mathbf{x})\neq(\Pi_{\tau_{0}}^{1}u)(\mathbf{x})$ in general in this case. To absorb the effect of $(\Pi^{*}u)(\mathbf{x})\neq(\Pi_{\tau_{0}}^{1}u)(\mathbf{x})$, we use the shape regularity condition (2.5) for $\tau_{0}$ and the Sobolev embedding inequality (2.7) for $T_{k}$. Therefore, we require that $\tau_{0}$ satisfies the minimum angle condition.
If $\mathbf{x}\in v(T_{k})$, in contrast, we have $(\Pi^{*}u)(\mathbf{x})=(\Pi_{T_{k}}^{1}u)(\mathbf{x})=(\Pi_{\tau_{0}}^{1}u)(\mathbf{x})$, and we do not need to impose such a condition (see Figure 4 (right)).
[…]
Figure  4. Two situations regarding Assumption 3 (3).
Suppose now that $v(\tau_{1})\cap v(\tau_{2})\neq\emptyset$, $\tau_{1}\in Q_{k}$, $\tau_{2}\in Q_{\ell}$, ($1\leq k<\ell\leq m$) and $\mathbf{x}\in v(\tau_{1})\cap v(\tau_{2})$ is an inner point of $\Omega$. If $(\Pi_{T_{k}}^{1}u)(\mathbf{x})\neq(\Pi_{T_{\ell}}^{1}u)(\mathbf{x})$, then $\Pi^{*}u$ is not well-defined. This case may occur if $\mathbf{x}$, $T_{k}$, and $T_{\ell}$ are in the situation depicted in Figure 5. Assumption 3 (4) is imposed to avoid such a situation. Three situations that satisfy Assumption 3 (4) are depicted in Figure 6.
[…]
Figure  5. A situation that makes $\Pi^{*}u$ be not well-defined.
[…]
Figure  6. Three situations that satisfy Assumption 3(4).
Suppose that $\tau_{1}\in Q_{k}$ ($1\leq k\leq m$) with $v(\tau_{1})\cap\partial\Omega\neq\emptyset$. The interpolation $\Pi^{*}u$ should then satisfy the boundary condition $\Pi^{*}u=0$ on $\partial\Omega$. Assumption 3(5) is imposed to ensure that this condition is fulfilled. Three situations that satisfy Assumption 3(5) are depicted in Figure 7.
[…]
Figure  7. Three situations that satisfy Assumption 3(5).
Theorem 6.
Suppose that the mesh $\mathcal{T}_{h}$ satisfies Assumption 3. The interpolation $\Pi^{*}u$ is then well defined for $u\in H^{2}(\Omega)$, and we have
(3.1) $|u-\Pi^{*}u|_{H^{1}(\Omega)}\leq Eh|u|_{H^{2}(\Omega)},\qquad\forall u\in H^{2}(\Omega),$
(3.2) $E:=\left(A_{\theta,\psi}^{2}(n+2+C^{2}M)+\frac{2(n+1)(n+2)\pi B_{\psi}^{2}C^{4-n}D_{{\theta}}MN}{n{\theta}}\right)^{1/2}.$
Proof.
Define $W\subset Q_{0}$ by $W:=\{\tau\in Q_{0}\mid\theta_{\tau}^{min}\geq{\theta}\}$. We decompose $|u-\Pi^{*}u|_{H^{1}(\Omega)}^{{2}}$ as
$|u-\Pi^{*}u|_{H^{1}(\Omega)}^{2}=\sum_{\tau\in W}|u-\Pi^{*}u|_{H^{1}(\tau)}^{2}$ $+\sum_{\tau\in Q_{0}\backslash W}|u-\Pi^{*}u|_{H^{1}(\tau)}^{2}$
$+\sum_{k=1}^{m}\sum_{\tau\in Q_{k}}|u-\Pi^{*}u|_{H^{1}(\tau)}^{2}.$
We estimate each term on the right-hand side. As
$\Pi^{*}u=\Pi_{\tau}^{1}u-(\Pi_{\tau}^{1}u-\Pi^{*}u)=\Pi_{\tau}^{1}u-\sum_{\mathbf{x}_{i}\in v(\tau)}(u(\mathbf{x}_{i})-w_{i})\varphi_{i}$
for $\tau\in W$, we have
$u-\Pi^{*}u=u-\Pi_{\tau}^{1}u+\sum_{\mathbf{x}_{i}\in v(\tau)}(u(\mathbf{x}_{i})-w_{i})\varphi_{i}.$
As $u(\mathbf{x}_{i})\neq w_{i}$ occurs to at most $n+1$ vertices of $\tau$, by the triangle and Cauchy–Schwarz inequalities, we have
$\left|u-\Pi^{*}u\right|_{H^{1}(\tau)}^{2}$ $\leq(n+2)\left|u-\Pi_{\tau}^{1}u\right|_{H^{1}(\tau)}^{2}+(n+2)\sum_{\mathbf{x}_{i}\in v(\tau)}\left|u(\mathbf{x}_{i})-w_{i}\right|^{2}|\varphi_{i}|_{H^{1}(\tau)}^{2}.$
Hence, we have
$\sum_{\tau\in W}\left|u-\Pi^{*}u\right|_{H^{1}(\tau)}^{2}$ $\leq(n+2)\sum_{\tau\in W}\left|u-\Pi_{\tau}u\right|_{H^{1}(\tau)}^{2}$
$\qquad+(n+2)\sum_{\tau\in W}\sum_{\mathbf{x}_{i}\in v(\tau)}\left|u(\mathbf{x}_{i})-w_{i}\right|^{2}|\varphi_{i}|_{H^{1}(\tau)}^{2}$
[…]
$\qquad+(n+2)\sum_{\tau\in W}\sum_{\mathbf{x}_{i}\in v(\tau)}\left|u(\mathbf{x}_{i})-w_{i}\right|^{2}|\varphi_{i}|_{H^{1}(\tau)}^{2}$
$\leq(n+2)A_{\theta,\psi}^{2}h^{2}\sum_{\tau\in W}|u|_{H^{2}(\tau)}^{2}$
$\qquad+(n+2)\sum_{\tau\in W}\frac{D_{{\theta}}h_{\tau}^{n-2}}{n(n-1)}\sum_{\mathbf{x}_{i}\in v(\tau)}\left|u(\mathbf{x}_{i})-w_{i}\right|^{2}.$
Here, we used the fact that
$|\nabla\varphi_{i}|$ $\leq\frac{1}{e_{\tau}}=\frac{|{L}_{\tau}|}{n|\tau|}\leq{\frac{h_{\tau}^{n-1}}{n(n-1)|\tau|}},$
[…]
where ${L}_{\tau}$ is the longest edge or widest facet of $\tau$, and $e_{\tau}$ is the height of $\tau$ with respect to ${L}_{\tau}$.
For the last term in the above inequality, we note that
$(n+2)\sum_{\tau\in W}$ $\frac{D_{{\theta}}h_{\tau}^{n-2}}{n(n-1)}\sum_{\mathbf{x}_{i}\in v(\tau)}|u(\mathbf{x}_{i})-w_{i}|^{2}$
$\leq(n+2)\frac{D_{{\theta}}h^{n-2}}{n(n-1)}\sum_{\tau\in W}\sum_{k=1}^{m}\sum_{\beta\in Q_{k}}\sum_{\mathbf{x}_{i}\in v(\tau)\cap v(\beta)}|u(\mathbf{x}_{i})-w_{i}|^{2}$
$\leq(n+2)\frac{D_{{\theta}}h^{n-2}}{n(n-1)}\sum_{\tau\in W}\sum_{k=1}^{m}\sum_{\beta\in Q_{k}}\sum_{\mathbf{x}_{i}\in v(\tau)\cap v(\beta)}\|u-\Pi_{T_{k}}^{1}u\|_{L^{\infty}(T_{k})}^{{2}}$
(by (2.7)) $\leq(n+2)\frac{D_{{\theta}}h^{n-2}}{n(n-1)}\sum_{\tau\in W}\sum_{k=1}^{m}\sum_{\beta\in Q_{k}}\sum_{\mathbf{x}_{i}\in v(\tau)\cap v(\beta)}B_{\psi}^{2}h_{T_{k}}^{4-n}|u|_{H^{2}(T_{k})}^{2}$
(by Assumption 3(1)) $\leq(n+2)\frac{D_{{\theta}}h^{n-2}}{n(n-1)}\sum_{\tau\in W}\sum_{k=1}^{m}\sum_{\beta\in Q_{k}}\sum_{\mathbf{x}_{i}\in v(\tau)\cap v(\beta)}B_{\psi}^{2}C^{4-n}h^{4-n}|u|_{H^{2}(T_{k})}^{2}$
$=(n+2)\frac{B_{\psi}^{2}C^{4-n}D_{{\theta}}h^{2}}{n(n-1)}\sum_{k=1}^{m}|u|_{H^{2}(T_{k})}^{2}\sum_{\beta\in Q_{k}}\sum_{\tau\in W}\sum_{\mathbf{x}_{i}\in v(\tau)\cap v(\beta)}1.$
Note that, for $\beta\in Q_{k}$ and each vertex $\mathbf{x}_{i}\in v(\beta)$, there are at most
(3.4) $\left\lfloor\frac{2(n-1)\pi}{{\theta}}\right\rfloor$
simplices such that $\tau\in W$ and $\mathbf{x}_{i}\in v(\tau)\cap v(\beta)$ because of Assumption 3(4). Here, $\lfloor x\rfloor$ is the floor function for $x\in\mathbb{R}$. Therefore, we have
$(n+2)$ $\frac{B_{\psi}^{2}C^{4-n}D_{{\theta}}h^{2}}{n(n-1)}\sum_{k=1}^{m}|u|_{H^{2}(T_{k})}^{2}\sum_{\beta\in Q_{k}}\sum_{\tau\in W}\sum_{\mathbf{x}_{i}\in v(\tau)\cap v(\beta)}1$
$\leq(n+2)\frac{B_{\psi}^{2}C^{4-n}D_{{\theta}}h^{2}}{n(n-1)}\sum_{k=1}^{m}|u|_{H^{2}(T_{k})}^{2}\sum_{\beta\in Q_{k}}(n+1)\left\lfloor\frac{2(n-1)\pi}{{\theta}}\right\rfloor$
(by Assumption 3(2)) $\leq\frac{2(n+1)(n+2)\pi B_{\psi}^{2}C^{4-n}D_{{\theta}}Nh^{2}}{n{\theta}}\sum_{k=1}^{m}|u|_{H^{2}(T_{k})}^{2}$
(by Assumption 3(1)) $\leq\frac{2(n+1)(n+2)\pi B_{\psi}^{2}C^{4-n}D_{{\theta}}MNh^{2}}{n{\theta}}|u|_{H^{2}(\Omega)}^{2}.$
Combining these inequalities, we obtain
$\sum_{\tau\in W}\left|u-\Pi^{*}u\right|_{H^{1}(\tau)}^{2}$ $\leq(n+2)A_{\theta,\psi}^{2}h^{2}\sum_{\tau\in W}|u|_{H^{2}(\tau)}^{2}$
$\quad+\frac{2(n+1)(n+2)\pi B_{\psi}^{2}C^{4-n}D_{{\theta}}MNh^{2}}{n{\theta}}|u|_{H^{2}(\Omega)}^{2}.$
Suppose now that $\tau\in Q_{0}\backslash W$. By Assumption 3(3), we have $\Pi^{*}u=\Pi_{\tau}^{1}u$ on $\tau$ as is pointed out in Remark 5. Hence, we have
$\sum_{\tau\in Q_{0}\backslash W}|u-\Pi^{*}u|_{H^{1}(\tau)}^{2}=\sum_{\tau\in Q_{0}\backslash W}|u-\Pi_{\tau}^{1}u|_{H^{1}(\tau)}^{2}\leq A_{\theta,\psi}^{2}h^{2}\sum_{\tau\in Q_{0}\backslash W}|u|_{H^{2}(\tau)}^{2}$
by (2.6).
As $\Pi^{*}u=\Pi_{T_{k}}^{1}u$ on $\tau\in Q_{k}$, $k=1,\dots,m$ by definition, we have
$\sum_{k=1}^{m}\sum_{\tau\in Q_{k}}|u-\Pi^{*}u|_{H^{1}(\tau)}^{2}$ $\leq\sum_{k=1}^{m}|u-\Pi_{T_{k}}^{1}u|_{H^{1}(T_{k})}^{2}$
(by (2.6)) $\leq\sum_{k=1}^{m}A_{\theta,\psi}^{2}h_{T_{k}}^{2}|u|_{H^{2}(T_{k})}^{2}$
(by Assumption 3(1)) $\leq A_{\theta,\psi}^{2}C^{2}Mh^{2}|u|_{H^{2}(\Omega)}^{2}.$
Gathering these three inequalities, we obtain
$|u-\Pi^{*}u|_{H^{1}(\Omega)}^{2}$ $=\sum_{\tau\in W}|u-\Pi^{*}u|_{H^{1}(\tau)}^{2}+\sum_{\tau\in Q_{0}\backslash W}|u-\Pi^{*}u|_{H^{1}(\tau)}^{2}$
$\hskip 96.73918pt+\sum_{k=1}^{m}\sum_{\tau\in Q_{k}}|u-\Pi^{*}u|_{H^{1}(\tau)}^{2}$
$\leq(n+2)A_{\theta,\psi}^{2}h^{2}\sum_{\tau\in W}|u|_{H^{2}(\tau)}^{2}$
$\quad+\frac{2(n+1)(n+2)\pi B_{\psi}^{2}C^{4-n}D_{{\theta}}MNh^{2}}{n{\theta}}|u|_{H^{2}(\Omega)}^{2}$
$\quad+A_{\theta,\psi}^{2}h^{2}\sum_{\tau\in Q_{0}\backslash W}|u|_{H^{2}(\tau)}^{2}+A_{\theta,\psi}^{2}C^{2}Mh^{2}|u|_{H^{2}(\Omega)}^{2}$
$\leq\left(A_{\theta,\psi}^{2}(n+2+C^{2}M)+\frac{2(n+1)(n+2)\pi B_{\psi}^{2}C^{4-n}D_{{\theta}}MN}{n{\theta}}\right)h^{2}|u|_{H^{2}(\Omega)},$
and Theorem 6 is proved. ∎
4. Simpler formulation for the case $n=2$.
For the case $n=2$, we introduce an assumption simpler than Assumption 3 that is still practical and applicable to many meshes.
Assumption 7.
$(1)$ There exist similar isosceles triangles $T_{1},\dots,T_{m}$ whose inner angles are ${\phi}$, ${\phi}$, and $\pi-2{\phi}$. We suppose that $\ring{T}_{k}\subset\Omega$, $k=1,\dots,m$, $\ring{T}_{i}\cap\ring{T}_{j}=\emptyset$ ($i\neq j$, $i,j=1,\dots,m$).
$(2)$ If $\theta_{\tau}^{max}>\pi-{\phi}$ for $\tau\in\mathcal{T}_{h}$, then $\ring{\tau}\subset\ring{T_{k}}$ with some $k$, and the longest edges of $\tau$ and $T_{k}$ coincide. Let $\mathbf{x}\in v(\tau)\backslash v(T_{k})$. All $\tau\in\mathcal{T}_{h}$ that share $\mathbf{x}$ must satisfy $\theta_{\tau}^{min}\geq{\phi}/2$.
[…]
Figure  8. A bad element $\tau=\triangle\mathbf{x}_{1}\mathbf{x}_{2}\mathbf{x}_{3}$ is covered virtually by a triangle $T_{k}$. They share the edge $\mathbf{x}_{1}\mathbf{x}_{2}$. The minimum inner angles of all elements except $\tau$ that share $\mathbf{x}_{3}$ are greater than or equal to ${\phi}/2$.
Figure 8 depicts a typical situation. Assumption 7 is a special case of Assumption 3. In fact, if Assumption 7 is satisfied, Assumption 3 is satisfied with $M=N=1$, $C=1$, $\theta=\phi/2$, and $\psi={\phi}$. Hence, we have the following corollary.
Corollary 8.
Suppose that the mesh $\mathcal{T}_{h}$ satisfies Assumption 7. Then, the interpolation $\Pi^{*}u$ is well defined for $u\in H^{2}(\Omega)$, and we have
$|u-\Pi^{*}u|_{H^{1}(\Omega)}\leq Eh|u|_{H^{2}(\Omega)},\qquad\forall u\in H^{2}(\Omega),$
where $E$ is a constant that depends only on ${\phi}$ and is independent of $h>0$.
Figure 9 shows a mesh that satisfies Assumption 3 but does not satisfy Assumption 7.
Figure  9. Bad elements that satisfy Assumption 3 but do not satisfy Assumption 7.
5. Comparison of the results with the results of prior works
In this section, we compare our approach with approaches adopted in prior works. As stated, the approach adopted in this paper was initiated by Kucěra [17] and Duprez, Lleras, and Lozinski [8].
First, we explain the approach of Kucěra [17]. Suppose that there exists a bad triangle $\tau$ with nodes $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, and $\mathbf{x}_{3}$ in the mesh at which the angle $\angle\mathbf{x}_{1}\mathbf{x}_{3}\mathbf{x}_{2}$ is close to $\pi$ (Figure 10 (left)). The interpolation $\Pi^{**}u$ for $u\in W^{2,\infty}(\tau)$ is then defined by
$(\Pi^{**}u)(\mathbf{x}_{1})=u(\mathbf{x}_{1}),\quad(\Pi^{**}u)(\mathbf{x}_{2})=u(\mathbf{x}_{2}),\quad(\nabla(\Pi^{**}u)(\mathbf{x}_{0}),\mathbf{v})=(\nabla u(\mathbf{x}_{0}),\mathbf{v}),$
where $\mathbf{x}_{0}$ is the foot on the edge $\overline{\mathbf{x}_{1}\mathbf{x}_{2}}$ from $\mathbf{x}_{3}$, and $\mathbf{v}:=(\mathbf{x}_{3}-\mathbf{x}_{0})/|\mathbf{x}_{3}-\mathbf{x}_{0}|$ (Figure 10 (right)).
$\mathbf{x}_{1}$$\mathbf{x}_{2}$$\mathbf{x}_{3}$
$\mathbf{x}_{1}$$\mathbf{x}_{2}$$\mathbf{x}_{3}$$\mathbf{x}_{0}$$\mathbf{v}$
Figure  10. A bad element and the modified Lagrange interpolation.
Let $\tau_{1}$ be another element that shares the node $\mathbf{x}_{3}$. Obviously, $(\Pi_{\tau_{1}}^{1}u)(\mathbf{x}_{3})\neq(\Pi^{**}u)(\mathbf{x}_{3})$ in general. To adjust the difference, the correction function $w$ is introduced so that
$(u+w)(\mathbf{x}_{3})=(\Pi^{**}u)(\mathbf{x}_{3}),\qquad\mathrm{supp}\,w=B(\mathbf{x}_{3},r_{\tau}),$
$B(\mathbf{x}_{3},r_{\tau}):=\{\mathbf{x}:|\mathbf{x}-\mathbf{x}_{3}|\leq r_{\tau}\},\quad r_{\tau}:=\frac{1}{2}\min\{|\mathbf{x}_{3}-\mathbf{x}_{1}|,|\mathbf{x}_{3}-\mathbf{x}_{2}|\}.$
Then, $\Pi^{**}u$ is defined by $\Pi_{\tau_{1}}^{1}(u+w)$ on $\tau_{1}$. To make $\Pi^{**}u$ well-defined globally, the supports of the correction functions should be disjoint (see Figure 11).
$\mathbf{x}_{1}$$\mathbf{x}_{2}$$\mathbf{x}_{3}$$\mathrm{supp}\,w$$\mathbf{x}_{4}$$\mathbf{x}_{5}$$\mathrm{supp}\,w$
Figure  11. The supports of the correction functions should be disjount.
An advantage of the approach adopted in [17] is that the surrounding elements of a bad element need to satisfy only the maximum angle condition, whereas Theorem 6 requires that such surrounding elements satisfy the minimum angle condition (see Assumption 3(3)). Hence, [17, Theorem 5] can deal with the mesh depicted in Figure 12, whereas Theorem 6 is not applied in this case.
Figure  12. A bad element is surrounded by elements that satisfy the maximum angle condition.
In contrast, the condition that the supports of the correction functions are disjoint is rather restrictive. For example, [17, Theorem 5] cannot deal with the bad elements depicted in Figure 6 (center), whereas Theorem 6 can be applied. In [17, Section 3.4], Kučera presented a way to deal with clusters of bad elements. However, each such cluster must be sufficiently far from other clusters and the boundary $\partial\Omega$. Note that Theorem 6 can be applied in the cases of a clusters of bad elements that are close or connected to other clusters or the boundary $\partial\Omega$, if Assumption 3(3) is satisfied.
Moreover, the approach in [17] deals with only triangular elements ($n=2$) in the $W^{2,\infty}$ setting, whereas Theorem 6 is applicable to the more general case of $n$-simplices ($n=2,3$) in the $H^{2}$ setting.
Second, we explain the approach of Duprez, Lleras, and Lozinski [8], which is slightly different from that of [17]. Their assumption is that each cluster of simplices that contains bad elements should be star-shaped and completely surrounded by good simplices.
Figure  13. A bed element is surrounded by good elements.
The interpolation $\Pi^{***}u$ is then defined by Taylor polynomials on the cluster. On the surrounding good elements, the Lagrange interpolation $\Pi_{\tau}^{1}u$ is modified to ensure the continuity of $\Pi^{***}u$ between elements. To evaluate the influence of the procedure, Duprez, Lleras, and Lozinski used $L^{\infty}$ error estimates of the Sobolev embedding, which is quite similar to our approach. Their main theorem [8, Proposition 2.7] can deal with both two- and three-dimensional cases in the $H^{2}$ setting.
Although [8, Proposition 2.7] is applicable to a wide range of configurations of bad elements, it seems that Theorem 6 can also deal with most of such cases. Furthermore, as Assumption 3 does not require that each cluster of bad elements is surrounded by good elements, Theorem 6 is applicable to a wider range of configurations than [8, Proposition 2.7]. For example, the mesh given in the numerical experiment (Figure 14) cannot be dealt with using the cited approach.
6. Numerical experiment
We present a numerical experiment to confirm the theoretical results obtained in the previous sections. Let $\Omega:=(0,1)\times(0,1)$. We consider the Poisson equation
(6.1) $-\Delta u=2(x(1-x)+y(1-y))\;\text{ in }\;\Omega,\qquad u=0\;\text{ on }\;\partial\Omega,$
whose exact solution is $u=x(1-x)y(1-y)$.
Let $K$ be a positive integer and $k=1/K$. To define the triangulation $\mathcal{T}_{h}$, we first divide $\Omega$ into $K^{2}$ small squares with side lengths $k$. Each small square is further subdivided into six triangles as depicted in Figure 14. Here, the coordinates of the points in each small square are $(x,y)$, $(x+k,y)$, $(x+k,y+k)$, $(x,y+k)$, $(x+k/2,y+\alpha k)$, and $(x+k/2,y+(1-\alpha)k)$, where $0<\alpha\ll 1$. Note that the smallest triangle becomes very bad when $\alpha$ approaches to zero. It is clear that this triangulation satisfies Assumption 7, and Corollary 8 is applicable to this triangulation $\mathcal{T}_{h}$.
$(x,y)$$(x+k,y)$$(x+k,y+k)$$(x,y+k)$$(x+\frac{k}{2},y+\alpha k)$$(x+\frac{k}{2},y+(1-\alpha)k)$
Figure  14. Triangulation $\mathcal{T}_{h}$ of $\Omega$ (left), and the subdivision of each small square (right).
We compute the piecewise linear finite element solutions $u_{h}\in S_{h0}$ of (6.1) on $\mathcal{T}_{h}$, and the error $|u-u_{h}|_{H^{1}(\Omega)}$ with varying $K$ and $\alpha$. Note that $h=(0.25+(1-\alpha)^{2})^{1/2}k$ if $\alpha<(2-\sqrt{3})/2=0.1339\dots$.
Table 1 presents the results of the numerical experiment. Here, $e_{h}$ stands for $|u-u_{h}|_{H^{1}(\Omega)}$. Although the meshes $\mathcal{T}_{h}$ contain many very bad elements, we observe that the error $|u-u_{h}|_{H^{1}(\Omega)}$ is of $\mathcal{O}(h)$ and consistent with Theorem 6 and Corollary 8.
Table 1. Results of the numerical experiment. In the table, $e_{h}$ stands for $|u-u_{h}|_{H^{1}(\Omega)}$.
$\alpha=0.1$ $\alpha=0.01$ $\alpha=0.0001$
$K$ $e_{h}$ $e_{h}/h$ $e_{h}$ $e_{h}/h$ $e_{h}$ $e_{h}/h$
10 1.8002 e-2 0.17485 2.0839 e-2 0.18789 2.1237 e-2 0.18997
20 9.0151 e-3 0.17512 1.0440 e-2 0.18827 1.0641 e-2 0.19036
40 4.5093 e-3 0.17519 5.2229 e-3 0.18836 5.3231 e-3 0.19046
80 2.2548 e-3 0.17521 2.6118 e-3 0.18839 2.6619 e-3 0.19049
160 1.1274 e-3 0.17521 1.3059 e-3 0.18866 1.3310 e-3 0.19049
7. Concluding remarks and future works
In this paper, we examined the error estimation of finite element solutions on meshes that contain many bad elements. Our research was motivated by the works of Kučera [17] and Duprez, Lleras, and Lozinski [8]. We showed that even if $\mathcal{T}_{h}$ contains many bad elements, finite element solutions $\{u_{h}\}$ converge to the exact solution $u\in H^{2}(\Omega)$ with order $\mathcal{O}(h)$, if each bad element is covered virtually by a simplex that satisfies the minimum angle condition. To be precise, we presented Assumption 3 that is very general and covers a wide range of triangulation patterns.
For the case $n=2$, we gave Assumption 7, which is simpler than Assumption 3 but is still practical and applicable to many meshes that cannot be dealt with using the methods of [17, 8].
In our numerical experiment, the finite element solutions behaved exactly as predicted from Theorem 6 and Corollary 8.
The results of this paper can be extended in many ways. An immediate extension may be the consideration of finite element solutions using higher order elements, and the authors are now working on this extension.